### v4

# GVE-LPA: Fast Label Propagation Algorithm (LPA) for Community Detection in Shared Memory Setting

Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for this purpose are crucial in various applications, particularly as datasets grow to substantial scales. This technical report presents an optimized parallel implementation of the Label Propagation Algorithm (LPA), a high speed community detection method, for shared memory multicore systems. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our LPA, which we term as GVE-LPA, outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively - achieving a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. In addition, GVE-LPA scales at a rate of 1.7×1.7$\times$1.7 × every doubling of threads.
Community detection, Parallel Label Propagation Algorithm (LPA)
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Community detection is the problem of identifying cohesive groups of vertices, also known as clusters, in complex networks that minimize cuts and maximize internal density (Garza and Schaeffer, 2019). This NP-hard problem finds applications in various domains, including drug discovery, protein annotation, topic exploration, anomaly detection, and criminal identification. Communities identified are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community (Gregory, 2010). The Label Propagation Algorithm (LPA), also known as RAK, (Blondel et al., 2008) is a popular heuristic-based approach for community detection, with the modularity metric (Newman, 2006) being used to measure the quality of communities identified.
In the past few years, there has been an unprecedented surge in the gathering of data and the depiction of their interconnections through graphs. This surge has underscored the need for devising efficient parallel algorithms tailored for identifying communities within massive networks. The significance of the multicore/shared memory environment in this context is paramount, owing to its energy-efficient nature and the prevalence of hardware equipped with substantial DRAM capacities. Existing studies on LPA propose algorithmic optimizations and parallelization techniques, but do not study efficient data structures for picking the most weighted label, to the best of our knowledge. Moreover, the proposed techniques are scattered over a number of papers.
1.1. Our Contributions
This report introduces GVE-LPA [^1] https://github.com/puzzlef/rak-communities-openmp, an optimized parallel implementation of LPA for shared memory multicores. On a machine with two 16-core Intel Xeon Gold 6226R processors, GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively. GVE-LPA is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, and achieves a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. With doubling of threads, GVE-LPA exhibits an average performance scaling of 1.7×1.7$\times$1.7 ×.
2. Related work
The Label Propagation Algorithm (LPA) is a diffusion-based technique for identifying communities. It is faster and more scalable than Louvain, as it does not require repeated optimization steps and is easier to parallelize (Newman, 2004; Raghavan et al., 2007). For parallelization of LPA, vertex assignment has been achieved with guided scheduling (Staudt and Meyerhenke, 2015), parallel bitonic sort (Soman and Narang, 2011), and pre-partitioning of the graph (Kuzmin et al., 2015).
In terms of quality, LPA may generate communities with low modularity, leading to the emergence of a dominant community that engulfs the majority of nodes, obstructing finer community structures (Gregory, 2010). To address instability of result, community structure quality, and performance due to vertex processing order, specific ordering strategies have been attempted. These include updating only active nodes (Xie and Szymanski, 2011) or prioritizing certain subsets based on properties like being core or boundary nodes (Gui et al., 2018). In some cases, oscillations in label assignments may occur instead of convergence. To address this, asynchronous mode (Leung et al., 2009), alternate updates of independent node subsets (Cordasco and Gargano, 2012), and parallel graph coloring techniques (Cordasco and Gargano, 2012) have been attempted. Raghavan et al. (Raghavan et al., 2007) report that after only five iterations, labels of $95\%$ of nodes converge to their final values.
Additional improvements upon the LPA include using a stable (non-random) mechanism of label choosing in the case of multiple best labels (Xing et al., 2014), addressing the issue of monster communities (Berahmand and Bouyer, 2018; Sattari and Zamanifar, 2018), identifying central nodes and combining communities for improved modularity (You et al., 2020), and using frontiers with alternating push-pull to reduce the number of edges visited and improve solution quality (Liu et al., 2020). A number of variants of LPA have been proposed, but the original formulation is still the simplest and most efficient (Garza and Schaeffer, 2019).
A few open source implementations and software packages have been developed for community detection using LPA. Fast Label Propagation Algorithm (FLPA) (Traag and Šubelj, 2023) is a fast variant of the LPA, which utilizes a queue-based approach to process only vertices with recently updated neighborhoods. NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It is implemented as a hybrid with C++ kernels and a Python frontend, and includes parallel implementation of LPA. igraph (Csardi et al., 2006) is a similar package, written in C, with Python, R, and Mathematica frontends. It is widely used in academic research, and includes implementation of LPA.
3. Preliminaries
Consider an undirected graph $G(V,E,w)$, where $V$ is the set of vertices, $E$ is the set of edges, and $w_{ij}=w_{ji}$ represents the weight associated with each edge. For unweighted graphs, we assume a unit weight for each edge, i.e., $w_{ij}=1$. The neighbors of a vertex $i$ are denoted as $J_{i}={j\ |\ (i,j)\in E}$, the weighted degree of each vertex as $K_{i}=\sum_{j\in J_{i}}w_{ij}$, the total number of vertices as $N=|V|$, the total number of edges as $M=|E|$, and the sum of edge weights in the undirected graph as $m=\sum_{i,j\in V}w_{ij}/2$.
3.1. Community detection
Disjoint community detection involves the identification of a community membership mapping, $C:V\rightarrow\Gamma$, where each vertex $i\in V$ is assigned a community ID $c\in\Gamma$, with $\Gamma$ representing the set of community IDs. We denote the vertices belonging to a community $c\in\Gamma$ as $V_{c}$, and the community to which a vertex $i$ belongs as $C_{i}$. Additionally, we denote the neighbors of vertex $i$ within community $c$ as $J_{i\rightarrow c}={j\ |\ j\in J_{i}\ and\ C_{j}=c}$, the sum of edge weights for those neighbors as $K_{i\rightarrow c}=\sum_{j\in J_{i\rightarrow c}}w_{ij}$, the edge weight within a community $c$ as $\sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=C_{j}=c}w_{ij}$, and the total edge weight of a community $c$ as $\Sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=c}w_{ij}$ (Leskovec, 2021).
3.2. Modularity
Modularity functions as a metric for evaluating the quality of communities identified by heuristic-based community detection algorithms. It is calculated as the difference between the fraction of edges within communities and the expected fraction under random edge distribution. It lies in a range of $[-0.5,1]$, where higher values signify superior quality (Brandes et al., 2007). The modularity $Q$ for the identified communities is computed using Equation 1, where $\delta$ denotes the Kronecker delta function ($\delta(x,y)=1$ if $x=y$, 00 otherwise).
(1) $Q=\frac{1}{2m}\sum_{(i,j)\in E}\left[w_{ij}-\frac{K_{i}K_{j}}{2m}\right]\delta(C_{i},C_{j})=\sum_{c\in\Gamma}\left[\frac{\sigma_{c}}{2m}-\left(\frac{\Sigma_{c}}{2m}\right)^{2}\right]$
3.3. Label Progagation Algorithm (LPA)
LPA (Raghavan et al., 2007) is a popular diffusion-based method for identifying communities of medium quality in large networks. It is simpler, faster, and more scalable compared to the Louvain method (Blondel et al., 2008). In LPA, each vertex $i$ is starts with a unique label (community id) $C_{i}$. During each iteration, vertices adopt the label with the highest interconnecting weight, as shown in Equation 2. Through this iterative process, a consensus among densely connected groups of vertices is achieved. The algorithm converges when at least $1-\tau$ fraction of vertices (where $\tau$ is the tolerance parameter) maintain their community membership. LPA exhibits a time complexity of $O(L|E|)$ and a space complexity of $O(|V|+|E|)$, with $L$ denoting the number of iterations performed (Raghavan et al., 2007). We have experimented with COPRA (Gregory, 2010), SLPA (Xie et al., 2011), and LabelRank (Xie et al., 2013), but found LPA to be the most performant, while yielding communities of equivalent quality (Sahu, 2023b).
(2) $C_{i}=\ \underset{c\ \in\ \Gamma}{\arg\max}{\sum_{j\in J_{i}\ |\ C_{j}=c}w_{ij}}$
4. Approach
4.1. Optimizations for LPA
We use a parallel implementation of LPA and experiment with different optimizations and parameter settings. We use the asynchronous version of LPA, wherein threads operate independently on distinct regions of the graph. This facilitates quicker convergence, but may introduce greater variability in the final result. Further, we observe that parallel LPA obtains communities of higher quality than its sequential implementation, possibly due to randomization. We allocate a separate hashtable per thread to keep track of the total weight of each unique label linked to a vertex.
Our optimizations include using OpenMP’s dynamic loop schedule, setting an initial tolerance of $0.05$, enabling vertex pruning, employing the strict version of LPA, and using fast collision-free per-thread hashtables which are well separated in their memory addresses (Far-KV). See below for the details on each optimization.
We evaluate multiple alternatives for each optimization, and show the relative time and the relative modularity of communities obtained by each alternative in Figure 3. We perform these tests on every graph in the dataset (refer to Table 1), conducting them five times on each graph to minimize the influence of noise. We then calculate the geometric mean for the runtime and arithmetic mean for the modularity, and represent them as ratios within each optimization category.
4.1.1. Adjusting OpenMP loop schedule
We attempt static, dynamic, guided, and auto loop scheduling approaches of OpenMP (each with a chunk size of $2048$) to parallelize LPA. We consider OpenMP’s dynamic loop schedule to be the best choice, due to its ability of work balancing among threads, and because it yields a runtime reduction of $27\%$ when compared to OpenMP’s auto loop schedule, while incurring only a $0.7\%$ reduction in the modularity of obtained communities (likely to be just noise).
4.1.2. Limiting the number of iterations
Restricting the number of iterations of LPA can ensure its termination within a reasonable number of iterations, but choosing a small limit may worsen the quality of communities obtained. Our results suggest that limiting the maximum number of iterations to $20$ strikes a good balance between runtime and modularity.
4.1.3. Adjusting tolerance
Using a small tolerance allows the algorithm to explore broader possibilities for community assignments, but comes at the cost of increased runtime. We find an initial tolerance of $0.05$ to be suitable. A tolerance of $0.1$ may also be acceptable, but provides a very small gain in performance when compared to a tolerance of $0.05$.
4.1.4. Vertex pruning
Vertex pruning is a method utilized to minimize unnecessary computation. In this approach, when a vertex alters its community, it assigns its neighbors for processing. Once a vertex has been processed, it is labeled as ineligible for further processing. However, this procedure incurs an additional overhead due to the marking and unmarking of vertices. Based on our findings, the employment of vertex pruning justifies this overhead and results in a performance enhancement of $17\%$. An illustration of vertex pruning optimization is shown in Figure 1.
Figure 1. Illustration of vertex pruning optimization: Once vertex $1$ is processed, it is unmarked. If vertex $1$ changes its community, its neighbors are marked for processing. Community membership of each vertex is indicated by border color, and marked vertices are highlighted in yellow (Sahu, 2023a).
4.1.5. Picking the best label
When there exist multiple labels connected to a vertex with maximum weight, we may randomly pick one of them (non-strict LPA), or pick only the first of them (strict LPA). We implement non-strict LPA using a simple modulo operator on the label id, as we observe that using xorshift based random number generator does not provide any advantage. Results indicate that the strict version of LPA is 1.5×1.5$\times$1.5 × faster than the non-strict approach, while also offering a gain in modularity of $2.1\%$.
4.1.6. Hashtable design
One can utilize C++’s inbuilt map as per-thread (independent) hashtables for the LPA algorithm. However, this exhibits poor performance. Therefore, we employ a key-list and a collision-free full-size values array to dramatically improve performance. However, if the memory addresses of the hashtables are nearby (Close-KV), even if each thread uses its own hashtable exclusively, the performance is not as high. This is possibly due to false cache-sharing. Alternatively, if we ensure that the memory address of each hashtable is farther away (Far-KV), the performance improves. Our results indicate that Far-KV has the best performance and is 15.8×15.8$\times$15.8 × times faster than Map, and 2.6×2.6$\times$2.6 × times faster than Close-KV with LPA. An illustration of Far-KV hashtable is in Figure 2.
Figure 2. Illustration of collision-free per-thread hashtables with well separated memory addresses (Far-KV) for two threads. Each hashtable comprises a keys vector, a values vector (of size $|V|$), and a key count ($N_{0}$/$N_{1}$). The value corresponding to each key is stored/accumulated at the index indicated by the key. As the key count of each hashtable is updated independently, we allocate it separately on the heap to avoid false cache sharing (Sahu, 2023a).
Figure 3. Impact of parameter controls and optimizations on the runtime and result quality (modularity) of LPA. We show the impact of each optimization upon the relative runtime on the left Y-axis, and upon the relative modularity on the right Y-axis.
4.2. Our optimized LPA implementation
We now explain the implementation of GVE-LPA in Algorithm 1. Its main step is the lpa() function (lines 7-13), which takes the input graph $G$ and assigns community memberships (or labels) $C$ to each vertex. In lines 8-9, we first initialize the labels $C$ of each vertex in $G$, and mark all vertices as unprocessed. We then perform iterations to propagate labels based on the weighted influence of neighboring vertices, limited to $MAX\_ITERATIONS$ (lines 10-12). In each iteration, we invoke the lpaMove() function to perform label propagation, and count the number of nodes with updated labels $\Delta N$ (line 11). If the ratio of $\Delta N$ to the total number of nodes $N$ is within a specified tolerance $\tau$, convergence has been achieved, and we terminate the loop (line 12). Upon completing all iterations, we return the final labels $C$ (line 13).
Algorithm 1 GVE-LPA: Our parallel LPA.
1:$G$: Input graph
2:$C$: Community membership of each vertex
3:$H_{t}$: Collision-free per-thread hashtable
4:$l_{i}$: Number of iterations performed
5:$\tau$: Per iteration tolerance
6:
7:function lpa($G$)
8:     Vertex membership: $C\leftarrow[0..|V|)$
9:     Mark all vertices in $G$ as unprocessed
[…]
11:         $\Delta N\leftarrow lpaMove(G,C)$
12:         if $\Delta N/N\leq\tau$ then break $\triangleright$ Converged?
13:     return $C$
14:
15:function lpaMove($G,C$)
16:     Changed vertices: $\Delta N\leftarrow 0$
17:     for all unprocessed $i\in V$ in parallel do
18:         Mark $i$ as processed (prune)
19:         $H_{t}\leftarrow scanCommunities(\{\},G,C,i)$
[…]
21:         $c^{*}\leftarrow$ Most weighted label to $i$ in $G$
22:         if $c^{*}=C[i]$ then continue
23:         $C[i]\leftarrow c^{*}$ ; $\Delta N\leftarrow\Delta N+1$
24:         Mark neighbors of $i$ as unprocessed
25:     return $\Delta N$
26:
27:function scanCommunities($H_{t},G,C,i$)
28:     for all $(j,w)\in G.edges(i)$ do
29:         if $i\neq j$ then $H_{t}[C[j]]\leftarrow H_{t}[C[j]]+w$
30:     return $H_{t}$
The lpaMove() function (lines 15-25) iterates over unprocessed vertices in parallel. For each unprocessed vertex $i$ in the graph $G$, we mark $i$ as processed - vertex pruning (line 18), obtain the total edge weight of connected labels in per-thread hashtable $H_{t}$ with the scanCommunities() function 19, and select the most weighted label $c^{*}$ (line 21). If $c^{*}$ is not the same as the current label of $i$, we update the label of $i$, increment the count of changed vertices $\Delta N$, and mark the neighbors of $i$ as unprocessed for the next iteration (lines 23-24). After having processed all vertices, we return the total number of vertices with updated labels $\Delta N$ (line 25). The scanCommunities() (lines 27-30) iterates over the neighbors of the current vertex $i$, excluding itself, and calculates the total weight of each label in the hashtable $H_{t}$.
5. Evaluation
5.1. Experimental Setup
We employ a server equipped with two Intel Xeon Gold 6226R processors, each featuring $16$ cores running at a clock speed of $2.90$ GHz. Every core is equipped with a $1$ MB L1 cache, a $16$ MB L2 cache, and shares a $22$ MB L3 cache. The system has $376$ GB of memory and runs on CentOS Stream 8. We use GCC 8.5 and OpenMP 4.5. Table 1 displays the graphs utilized in our experiments, all sourced from the SuiteSparse Matrix Collection (Kolodziej et al., 2019).
Table 1. List of $13$ graphs obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with ***). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities identified with GVE-LPA.
Graph $|V|$ $|E|$ $D_{avg}$ $|\Gamma|$
Web Graphs (LAW)
$*{}^{*}$ 7.41M 341M 41.0 147K
$*{}^{*}$ 18.5M 567M 16.1 383K
$*{}^{*}$ 22.7M 1.21B 28.2 213K
$*{}^{*}$ 39.5M 1.73B 23.7 677K
$*{}^{*}$ 118M 1.89B 8.6 6.48M
$*{}^{*}$ 41.3M 2.19B 27.9 611K
$*{}^{*}$ 50.6M 3.80B 38.5 284K
Social Networks (SNAP)
com-LiveJournal 4.00M 69.4M 17.4 2.19K
com-Orkut 3.07M 234M 76.2 49
Road Networks (DIMACS10)
asia_osm 12.0M 25.4M 2.1 278K
europe_osm 50.9M 108M 2.1 1.52M
Protein k-mer Graphs (GenBank)
kmer_A2a 171M 361M 2.1 40.1M
kmer_V1r 214M 465M 2.2 48.7M
(a) Runtime in seconds (logarithmic scale) with FLPA, igraph LPA, NetworKit LPA, and GVE-LPA
(b) Speedup (logarithmic scale) of GVE-LPA with respect to FLPA, igraph LPA, NetworKit LPA.
(c) Modularity of communities obtained with FLPA, igraph LPA, NetworKit LPA, and GVE-LPA.
Figure 4. Runtime, speedup, and modularity of communities obtained with FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph. FLPA and igraph LPA fail to run on kmer_A2a and kmer_V1r graphs, and thus their results are not shown.
(a) Runtime in seconds (logarithmic scale) with GVE-Louvain and GVE-LPA
(b) Speedup of GVE-LPA with respect to GVE-Louvain.
(c) Modularity of communities obtained with GVE-Louvain and GVE-LPA.
Figure 5. Runtime in seconds (logarithmic scale), speedup, and modularity of communities obtained with GVE-Louvain (Sahu, 2023a) and GVE-LPA for each graph in the dataset.
5.2. Comparing Performance of GVE-LPA
We now compare the performance of GVE-LPA with igraph LPA (Csardi et al., 2006), FLPA (Traag and Šubelj, 2023), and NetworKit LPA (Staudt et al., 2016). igraph LPA and FLPA are both sequential implementations, while NetworKit LPA is parallel. For FLPA, we checkout the suitable branch with modified igraph_community_label_propagation() function, update the label propagation example in C to load the input graph from a file and measure runtime of igraph_community_label_propagation () invoked with IGRAPH_LPA_FAST variant, using gettimeofday(). For igraph LPA, we follow a similar process as FLPA, but checkout code from the master branch of igraph. For NetworKit, we employ a Python script to invoke PLP (Parallel Label Propagation), and measure the total time reported with getTiming(). For each graph, we measure the NetworKit LPA runtime five times for averaging. Due to time constraints and long running times of igraph LPA, we ran it only once per graph. Additionally, we record the modularity of communities obtained from each implementation.
Figure 4(a) shows the runtimes of FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph in the dataset, while Figure 4(b) shows the speedup of GVE-LPA with respect to each aforementioned implementation. igraph LPA fails to run on protein k-mer graphs, i.e., kmer_A2a and kmer_V1r, and hence its results on these graphs are not shown. GVE-LPA exhibits an average speedup of 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × over FLPA, igraph LPA, and NetworKit LPA respectively. On the sk-2005 graph, GVE-LPA identifies communities in $2.7$ seconds, achieving a processing rate of $1.4$ billion edges/s. Figure 4(c) shows the modularity of communities obtained by each implementation. On average, GVE-LPA obtains $6.6\%$ / $0.2\%$ higher modularity than FLPA (especially on road networks) and igraph LPA respectively, and $4.1\%$ lower modularity than NetworKit LPA (especially on protein k-mer graphs).
Figure 6. Runtime $/|E|$ factor with GVE-LPA for each graph in the dataset.
Figure 7. Average speedup of GVE-Louvain and GVE-LPA with increasing number of threads (in multiples of 2).
Next, we proceed to compare the performance of GVE-LPA with GVE-Louvain in Figure 5. Figures 5(a), 5(b), and 5(c) present the runtimes, speedup (of GVE-LPA with respect to GVE-Louvain), and modularity of GVE-Louvain and GVE-LPA for each graph in the dataset. GVE-LPA offers a mean speedup of 5.4×5.4$\times$5.4 × compared to GVE-Louvain, particularly on social networks, road networks, and protein k-mer graphs, while achieving on average $10.9\%$ lower modularity, especially on social networks and protein k-mer graphs.
In Figure 6, we assess the runtime per edge of GVE-LPA. Results indicate that GVE-LPA exhibits a higher $\text{runtime}/|E|$ factor on graphs with a low average degree (road networks and protein k-mer graphs) and on graphs with a poor community structure (com-LiveJournal and com-Orkut).
5.3. Strong Scaling of GVE-LPA
Finally, we assess the strong scaling performance of GVE-LPA. To this end, we vary the number of threads from $1$ to $64$ in multiples of $2$ for each input graph, and measure the time required for GVE-LPA to identify communities (averaged over five runs). Figure 7 shows the average strong-scaling of of GVE-LPA, in comparison with GVE-Louvain. With 32 threads, GVE-LPA achieves a speedup of 13.5×13.5$\times$13.5 × compared to single-threaded execution, indicating a performance increase of 1.7×1.7$\times$1.7 × for each doubling of threads. At 64 threads, GVE-LPA is moderately impacted by NUMA effects, and offers a speedup of 18.1×18.1$\times$18.1 ×. This contrasts with GVE-Louvain, which appears to be significantly affected by NUMA effects.
6. Conclusion
In summary, this study focuses on optimizing the Label Propagation Algorithm (LPA), a high-speed community detection algorithm, in the shared memory setting. We consider 6 different optimizations, which significantly improve the performance of the algorithm. Comparative assessments against competitive open-source implementations (FLPA) and packages (igraph and NetworKit) indicate that GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97,000×97,000$\times$97 , 000 ×, and 40×40$\times$40 × respectively, while identifying communities of $6.6\%$ / $0.2\%$ higher quality than FLPA and igraph LPA respectively, and of $4.1\%$ lower quality than NetworKit LPA.
On a web graph with $3.8$ billion edges, GVE-LPA identifies communities in $2.7$ seconds, and thus achieves a processing rate of $1.4$ billion edges/s. GVE-LPA is ideal for applications requiring a high-speed clustering algorithm while accepting a compromise in clustering quality, as it is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, has a strong scaling factor of 1.7×1.7$\times$1.7 × for every doubling of threads, but identifies communities with on average $10.9\%$ lower modularity than GVE-Louvain. Future research could explore dynamic algorithms for LPA to accommodate evolving graphs in real-world scenarios. This would allow interactive updation of community memberships of vertices on large graphs.
Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Balavarun Pedapudi, Souvik Karfa, and Vincent Traag for their support.