### v6

# GVE-LPA: Fast Label Propagation Algorithm (LPA) for Community Detection in Shared Memory Setting

Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for this purpose are crucial in various applications, particularly as datasets grow to substantial scales. This technical report presents an optimized parallel implementation of the Label Propagation Algorithm (LPA), a high speed community detection method, for shared memory multicore systems. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our LPA, which we term as GVE-LPA, outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97000×97000$\times$97000 ×, and 40×40$\times$40 × respectively - achieving a processing rate of $1.4B$ edges/s on a $3.8B$ edge graph. In addition, GVE-LPA scales at a rate of 1.7×1.7$\times$1.7 × every doubling of threads.
Community detection, Parallel Label Propagation Algorithm (LPA)
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Community detection is the problem of identifying cohesive groups of vertices, also known as clusters, in complex networks that minimize cuts and maximize internal density (Garza and Schaeffer, 2019). Communities identified are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community (Gregory, 2010). One of the difficulties in community detection is the lack of apriori knowledge on the number and size distribution of communities (Blondel et al., 2008).
Community detection finds applications in various domains. In drug discovery, it helps identify groups of similar compounds or target proteins, thereby facilitating the discovery of new therapeutic agents (Ma et al., 2019; Udrescu et al., 2020). In the health domain, it aids in understanding the dynamics of groups susceptible to epidemic diseases (Salathé and Jones, 2010), detecting diseases like lung cancer (Bechtel et al., 2005), and categorizing tumor types using genomic datasets (Haq and Wang, 2016). Additionally, it has been applied to elucidate the structure and evolution of metabolic networks (Pfeiffer et al., 2005; Kim et al., 2009), Gene Regulatory Networks (GRNs) (Rivera et al., 2010), and Lateral Gene Transfer (LGT) networks (Popa et al., 2011). Another significant application of community detection is in the analysis of human brain networks (Bullmore and Sporns, 2009; He and Evans, 2010). In ecological studies, it is used to determine if food webs are organized into compartments, where species within the same compartment frequently interact among themselves but have fewer interactions with species in different compartments (Krause et al., 2003; Rezende et al., 2009; Guimerà et al., 2010).
The Label Propagation Algorithm (LPA), also known as RAK, (Raghavan et al., 2007) is a popular diffusion-based approach for identifying communities. It is faster and more scalable than the Louvain method (Blondel et al., 2008), another high-quality community detection algorithm, as it does not require repeated optimization steps and is easier to parallelize (Newman, 2004; Raghavan et al., 2007). Due to this, LPA has been used in a number of other graph problems and applications, such as, finding connected components (Stergiou et al., 2018), graph partitioning (Slota et al., 2014; Wang et al., 2014; Meyerhenke et al., 2014, 2016, 2017; Bae et al., 2020; Akhremtsev et al., 2020; Slota et al., 2020; Zhang et al., 2020a), hypergraph partitioning (Henne, 2015; Gottesbüren et al., 2021), graph coarsening (Valejo et al., 2020), vertex reordering and graph compression (Boldi et al., 2011), unsupervised part-of-speech tagging (Das and Petrov, 2011), sectionalizing power systems (Aziz et al., 2023). Further, Peng et al. (Peng et al., 2014) observe that, while LPA achieves lower modularity scores (a metric for assessing community quality (Newman, 2004)) than the Louvain method — a lower modularity score for LPA implies that it may struggle to properly cluster nodes that are not well-connected to any particular community— it attains the highest Normalized Mutual Information (NMI) score when compared to the ground truth.
A large number of LPA variants have also been proposed (Li et al., 2015a; Farnadi et al., 2015; Li et al., 2015b; SHEN and YANG, 2016; Mohan et al., 2017; Zhang et al., 2017; Berahmand and Bouyer, 2018a; Ma et al., 2018; Sattari and Zamanifar, 2018b; Zheng et al., 2018; Xu et al., 2019; Zarei et al., 2020; Maleki et al., 2020; Zhang et al., 2020b; El Kouni et al., 2021; Roghani et al., 2021; Zhang et al., 2023). Existing studies on LPA propose a number of algorithmic optimizations and parallelization techniques, but there continue to be a number of issue with each implementation. Further, the race for AI development has caused the prices of GPUs to skyrocket. These issues, and the large applicability of LPA, motivates us to present GVE-LPA, [^1] https://github.com/puzzlef/rak-communities-openmp an efficient implementation of LPA for multicore architectures.
2. Related work
For parallelization of LPA, vertex assignment has been achieved with guided scheduling (Staudt and Meyerhenke, 2015), parallel bitonic sort (Soman and Narang, 2011), and pre-partitioning of the graph (Kuzmin et al., 2015). Additional improvements upon the LPA include using a stable (non-random) mechanism of label choosing in the case of multiple best labels (Xing et al., 2014), addressing the issue of monster communities (Berahmand and Bouyer, 2018b; Sattari and Zamanifar, 2018a), identifying central nodes and combining communities for improved modularity (You et al., 2020), and using frontiers with alternating push-pull to reduce the number of edges visited and improve solution quality (Liu et al., 2020).
A few open source implementations have been developed for community detection using LPA. Fast Label Propagation Algorithm (FLPA) (Traag and Šubelj, 2023) is a fast variant of the LPA, which utilizes a queue-based approach to process only vertices with recently updated neighborhoods. NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It is implemented as a hybrid with C++ kernels and a Python frontend, and includes parallel implementation of LPA. igraph (Csardi et al., 2006) is a similar package, written in C, with Python, R, and Mathematica frontends. It is widely used in research, and includes an implementation of LPA.
We now discuss the implementation of LPA in NetworKit, igraph, and FLPA. The parallel implementation of LPA in NetworKit is given in NetworKit::PLP::run(). This function starts with a unique label for each node. To check for convergence, they use a tolerance of $10^{-5}$, i.e., the algorithm converges once the labels of less than $0.001\%$ of vertices change. This is also referred to as threshold heuristic. To track active nodes, a boolean flag vector is employed, serving as a vertex pruning optimization. A parallel for loop, which uses OpenMP’s guided schedule, then processes only the active nodes. For storing label weights, an std::map is used for each vertex. In contrast, the LPA implementation of igraph and FLPA is sequential. igraph’s LPA implementation, found in the function igraph_community_label_propagation(), shuffles the node order in each iteration. It keeps track of the dominant labels for each new neighbor of a vertex and selects one of the dominant labels using a random number generator. The hashtable is cleared by iterating through the neighbors. Unlike NetworKit, igraph does not use the pruning optimization. It alternates between label updating and control iterations, checking if the current label of a node is not dominant during control iterations, with no tolerance for non-dominant labels. The FLPA implementation in igraph, detailed in the function igraph_i_community_label_propagation(), uses a dequeue for managing a set of unprocessed vertices, which indicates convergence when empty. FLPA does not shuffle for random node order. If a label change occurs, FLPA considers neighbors that are not in the same community.
However, we identify a few issues with the LPA implementations of NetworKit, igraph, and FLPA. In order to assign a unique label to each node in the graph, NetworKit uses OpenMP’s plain parallel for, i.e., it uses static loop scheduling with a chunk size of $1$. This can result in false sharing as threads make writes to consecutive locations in memory. Next, to keep track of the weights associated with each label for a given vertex, they use an std::map. Our experiments show that this is quite inefficient. We, in contrast, use a per-thread keys list and a per-thread full-size values array, as our hashtable. In order to check for convergence, NetworKit uses a tolerance of $10^{-5}$. However, we observe that a tolerance of $10^{-2}$ is generally obtains community of nearly the same quality (in terms of modularity), but converges much faster. Further, NetworKit uses atomic operations on a shared variable to keep a count of the number of updated vertices. This can introduce contention. We, instead, make use of parallel reduce for this. Finally, NetworKit uses a boolean vector to keep track of the set of active nodes. However, our experiments show that using an 8-bit integer flag vector is more efficient.
We now discuss issues in the LPA implementation of igraph. igraph randomly shuffles the processing order of nodes in each iteration. This can get quite expensive. igraph checks for convergence in alternate iterations, and algorithm is considered to have converged only if the labels of all nodes are dominant, i.e., are the most popular. Thus it can take a large number of iterations to converge, with minimal gain in community quality. In addition, igraph does not use vertex pruning optimization, i.e., it does not maintain the set of active nodes. Finally, given multiple dominant labels, igraph select a random dominant label as the label of the given node — however, random number generation is slow. We use the dominant label with the smallest ID as the label of the given vertex.
Finally, we discuss the issues in FLPA’s implementation. Similar to igraph LPA, given multiple dominant labels, FLPA selects a random dominant label as the label of the given node — again, random number generation is slow. Further, FLPA considers the algorithm to have converged only when there are no active nodes. Thus it can take a large number of iterations to converge, with minimal gain in community quality.
3. Preliminaries
Consider an undirected graph $G(V,E,w)$, where $V$ is the set of vertices, $E$ is the set of edges, and $w_{ij}=w_{ji}$ represents the weight associated with each edge. For unweighted graphs, we assume a unit weight for each edge, i.e., $w_{ij}=1$. The neighbors of a vertex $i$ are denoted as $J_{i}={j\ |\ (i,j)\in E}$, the weighted degree of each vertex as $K_{i}=\sum_{j\in J_{i}}w_{ij}$, the total number of vertices as $N=|V|$, the total number of edges as $M=|E|$, and the sum of edge weights in the undirected graph as $m=\sum_{i,j\in V}w_{ij}/2$.
3.1. Community detection
Disjoint community detection involves the identification of a community membership mapping, $C:V\rightarrow\Gamma$, where each vertex $i\in V$ is assigned a community ID $c\in\Gamma$, with $\Gamma$ representing the set of community IDs. We denote the vertices belonging to a community $c\in\Gamma$ as $V_{c}$, and the community to which a vertex $i$ belongs as $C_{i}$. Additionally, we denote the neighbors of vertex $i$ within community $c$ as $J_{i\rightarrow c}={j\ |\ j\in J_{i}\ and\ C_{j}=c}$, the sum of edge weights for those neighbors as $K_{i\rightarrow c}=\sum_{j\in J_{i\rightarrow c}}w_{ij}$, the edge weight within a community $c$ as $\sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=C_{j}=c}w_{ij}$, and the total edge weight of a community $c$ as $\Sigma_{c}=\sum_{(i,j)\in E\ and\ C_{i}=c}w_{ij}$ (Leskovec, 2021).
3.2. Modularity
Modularity functions as a metric for evaluating the quality of communities identified by heuristic-based community detection algorithms. It is calculated as the difference between the fraction of edges within communities and the expected fraction under random edge distribution. It lies in a range of $[-0.5,1]$, where higher values signify superior quality (Brandes et al., 2007). The modularity $Q$ for the identified communities is computed using Equation 1, where $\delta$ denotes the Kronecker delta function ($\delta(x,y)=1$ if $x=y$, 00 otherwise).
(1) $Q=\frac{1}{2m}\sum_{(i,j)\in E}\left[w_{ij}-\frac{K_{i}K_{j}}{2m}\right]\delta(C_{i},C_{j})=\sum_{c\in\Gamma}\left[\frac{\sigma_{c}}{2m}-\left(\frac{\Sigma_{c}}{2m}\right)^{2}\right]$
3.3. Label Progagation Algorithm (LPA)
LPA (Raghavan et al., 2007) is a popular diffusion-based method for identifying communities of medium quality in large networks. It is simpler, faster, and more scalable compared to the Louvain method (Blondel et al., 2008). In LPA, each vertex $i$ is starts with a unique label (community id) $C_{i}$. During each iteration, vertices adopt the label with the highest interconnecting weight, as shown in Equation 2. Through this iterative process, a consensus among densely connected groups of vertices is achieved. The algorithm converges when at least $1-\tau$ fraction of vertices (where $\tau$ is the tolerance parameter) maintain their community membership. LPA exhibits a time complexity of $O(L|E|)$ and a space complexity of $O(|V|+|E|)$, with $L$ denoting the number of iterations performed (Raghavan et al., 2007). We have experimented with COPRA (Gregory, 2010), SLPA (Xie et al., 2011), and LabelRank (Xie et al., 2013), but found LPA to be the most performant, while yielding communities of equivalent quality (Sahu, 2023b).
(2) $C_{i}=\ \underset{c\ \in\ \Gamma}{\arg\max}{\sum_{j\in J_{i}\ |\ C_{j}=c}w_{ij}}$
4. Approach
4.1. Optimizations for LPA
We use an asynchronous parallel implementation of LPA, wherein threads operate independently on distinct regions of the graph. This facilitates quicker convergence, but may introduce greater variability in the final result. Further, we observe that parallel LPA obtains communities of higher quality than its sequential implementation, possibly due to randomization. We allocate a separate hashtable per thread to keep track of the total weight of each unique label linked to a vertex.
We now evaluate alternatives for a few optimizations, on the system mentioned in Section 5.1, for every graph in the dataset (Table 1) — conducting them five times on each graph to minimize the influence of noise. We then calculate the geometric mean for the runtime and arithmetic mean for the modularity, and represent them as ratios within each optimization category, as shown in Figure 3. While none of our optimizations are particularly novel, we carefully address each one and combine them to achieve superior performance compared to other implementations of LPA.
4.1.1. Adjusting OpenMP loop schedule
We attempt various loop scheduling approaches of OpenMP — each using a chunk size of $2048$ (we observe this strikes a balance between minimizing load imbalance and reducing scheduling overhead) to parallelize LPA. We consider OpenMP’s dynamic loop schedule to be the best choice, due to its ability of work balancing among threads, and because it yields a runtime reduction of $27\%$ when compared to OpenMP’s auto loop schedule, while incurring only a $0.7\%$ reduction in the modularity of obtained communities (likely to be just noise).
4.1.2. Limiting the number of iterations
Restricting the number of iterations of LPA can ensure its termination within a reasonable number of iterations, but choosing a small limit may worsen the quality of communities obtained. Our results suggest that limiting the maximum number of iterations to $20$ strikes a good balance between runtime and modularity.
4.1.3. Adjusting tolerance
Using a small tolerance allows the algorithm to explore broader possibilities for community assignments, but comes at the cost of increased runtime. We find an initial tolerance of $0.05$ to be suitable. A tolerance of $0.1$ may also be acceptable, but provides a very small gain in performance when compared to a tolerance of $0.05$.
4.1.4. Vertex pruning
Vertex pruning is a method utilized to minimize unnecessary computation. In this approach, when a vertex alters its community, it assigns its neighbors for processing. Once a vertex has been processed, it is labeled as ineligible for further processing. However, this procedure incurs an additional overhead due to the marking and unmarking of vertices. Based on our findings, the employment of vertex pruning justifies this overhead and results in a performance enhancement of $17\%$. An illustration of vertex pruning optimization is shown in Figure 1.
Figure 1. Illustration of vertex pruning optimization: Once vertex $1$ is processed, it is unmarked. If vertex $1$ changes its community, its neighbors are marked for processing. Community membership of each vertex is indicated by border color, and marked vertices are highlighted in yellow (Sahu, 2023a).
4.1.5. Picking the best label
When there exist multiple labels connected to a vertex with maximum weight, we may randomly pick one of them (non-strict LPA), or pick only the first of them (strict LPA). We implement non-strict LPA using a simple modulo operator on the label id, as we observe that using xorshift based random number generator does not provide any advantage. Results indicate that the strict version of LPA is 1.5×1.5$\times$1.5 × faster than the non-strict approach, while also offering a gain in modularity of $2.1\%$.
4.1.6. Hashtable design
One can utilize C++’s inbuilt map as per-thread (independent) hashtables for the LPA algorithm. However, this exhibits poor performance. Therefore, we employ a key-list and a collision-free full-size values array to dramatically improve performance. However, if the memory addresses of the hashtables are nearby (Close-KV), even if each thread uses its own hashtable exclusively, the performance is not as high. This is possibly due to false cache-sharing. Alternatively, if we ensure that the memory address of each hashtable is farther away (Far-KV), the performance improves. Our results indicate that Far-KV has the best performance and is 15.8×15.8$\times$15.8 × times faster than Map, and 2.6×2.6$\times$2.6 × times faster than Close-KV with LPA. An illustration of Far-KV hashtable is in Figure 2.
Figure 2. Illustration of collision-free per-thread hashtables with well separated memory addresses (Far-KV) for two threads. Each hashtable comprises a keys vector, a values vector (of size $|V|$), and a key count ($N_{0}$/$N_{1}$). The value corresponding to each key is stored/accumulated at the index indicated by the key. As the key count of each hashtable is updated independently, we allocate it separately on the heap to avoid false cache sharing (Sahu, 2023a).
Figure 3. Impact of parameter controls and optimizations on the runtime and result quality (modularity) of LPA. We show the impact of each optimization upon the relative runtime on the left Y-axis, and upon the relative modularity on the right Y-axis.
4.2. Our optimized LPA implementation
We now explain the implementation of GVE-LPA in Algorithm 1. Its main step is the lpa() function (lines 7-13), which takes the input graph $G$ and assigns community memberships (or labels) $C$ to each vertex. In lines 8-9, we first initialize the labels $C$ of each vertex in $G$, and mark all vertices as unprocessed. We then perform iterations to propagate labels based on the weighted influence of neighboring vertices, limited to $MAX\_ITERATIONS$ (lines 10-12). In each iteration, we invoke the lpaMove() function to perform label propagation, and count the number of nodes with updated labels $\Delta N$ (line 11). If the ratio of $\Delta N$ to the total number of nodes $N$ is within a specified tolerance $\tau$, convergence has been achieved, and we terminate the loop (line 12). Upon completing all iterations, we return the final labels $C$ (line 13).
Algorithm 1 GVE-LPA: Our parallel LPA.
1:$G$: Input graph
2:$C$: Community membership of each vertex
3:$H_{t}$: Collision-free per-thread hashtable
4:$l_{i}$: Number of iterations performed
5:$\tau$: Per iteration tolerance
6:
7:function lpa($G$)
8:     Vertex membership: $C\leftarrow[0..|V|)$
9:     Mark all vertices in $G$ as unprocessed
[…]
11:         $\Delta N\leftarrow lpaMove(G,C)$
12:         if $\Delta N/N\leq\tau$ then break $\triangleright$ Converged?
13:     return $C$
14:
15:function lpaMove($G,C$)
16:     Changed vertices: $\Delta N\leftarrow 0$
17:     for all unprocessed $i\in V$ in parallel do
18:         Mark $i$ as processed (prune)
19:         $H_{t}\leftarrow scanCommunities(\{\},G,C,i)$
[…]
21:         $c^{*}\leftarrow$ Most weighted label to $i$ in $G$
22:         if $c^{*}=C[i]$ then continue
23:         $C[i]\leftarrow c^{*}$ ; $\Delta N\leftarrow\Delta N+1$
24:         Mark neighbors of $i$ as unprocessed
25:     return $\Delta N$
26:
27:function scanCommunities($H_{t},G,C,i$)
28:     for all $(j,w)\in G.edges(i)$ do
29:         if $i\neq j$ then $H_{t}[C[j]]\leftarrow H_{t}[C[j]]+w$
30:     return $H_{t}$
The lpaMove() function (lines 15-25) iterates over unprocessed vertices in parallel. For each unprocessed vertex $i$ in the graph $G$, we mark $i$ as processed - vertex pruning (line 18), obtain the total edge weight of connected labels in per-thread hashtable $H_{t}$ with the scanCommunities() function 19, and select the most weighted label $c^{*}$ (line 21). If $c^{*}$ is not the same as the current label of $i$, we update the label of $i$, increment the count of changed vertices $\Delta N$, and mark the neighbors of $i$ as unprocessed for the next iteration (lines 23-24). After having processed all vertices, we return the total number of vertices with updated labels $\Delta N$ (line 25). The scanCommunities() (lines 27-30) iterates over the neighbors of the current vertex $i$, excluding itself, and calculates the total weight of each label in the hashtable $H_{t}$.
5. Evaluation
5.1. Experimental Setup
We employ a server equipped with two Intel Xeon Gold 6226R processors, each featuring $16$ cores running at a clock speed of $2.90$ GHz. Every core is equipped with a $1$ MB L1 cache, a $16$ MB L2 cache, and shares a $22$ MB L3 cache. The system has $376$ GB of memory and runs on CentOS Stream 8. We use GCC 8.5 and OpenMP 4.5. Table 1 displays the graphs utilized in our experiments, all sourced from the SuiteSparse Matrix Collection (Kolodziej et al., 2019).
Table 1. List of $13$ graphs obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with *). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities identified with GVE-LPA.
Graph $|V|$ $|E|$ $D_{avg}$ $|\Gamma|$
Web Graphs (LAW)
indochina-2004 7.41M 341M 41.0 147K
uk-2002 18.5M 567M 16.1 383K
arabic-2005 22.7M 1.21B 28.2 213K
uk-2005 39.5M 1.73B 23.7 677K
webbase-2001 118M 1.89B 8.6 6.48M
it-2004 41.3M 2.19B 27.9 611K
sk-2005 50.6M 3.80B 38.5 284K
Social Networks (SNAP)
com-LiveJournal 4.00M 69.4M 17.4 2.19K
com-Orkut 3.07M 234M 76.2 49
Road Networks (DIMACS10)
asia_osm 12.0M 25.4M 2.1 278K
europe_osm 50.9M 108M 2.1 1.52M
Protein k-mer Graphs (GenBank)
kmer_A2a 171M 361M 2.1 40.1M
kmer_V1r 214M 465M 2.2 48.7M
(a) Runtime in seconds (logarithmic scale) with FLPA, igraph LPA, NetworKit LPA, and GVE-LPA
(b) Speedup (logarithmic scale) of GVE-LPA with respect to FLPA, igraph LPA, NetworKit LPA.
(c) Modularity of communities obtained with FLPA, igraph LPA, NetworKit LPA, and GVE-LPA.
Figure 4. Runtime, speedup, and modularity of communities obtained with FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph. FLPA and igraph LPA fail to run on kmer_A2a and kmer_V1r graphs, and thus their results are not shown.
(a) Runtime in seconds (logarithmic scale) with GVE-Louvain and GVE-LPA
(b) Speedup of GVE-LPA with respect to GVE-Louvain.
(c) Modularity of communities obtained with GVE-Louvain and GVE-LPA.
Figure 5. Runtime in seconds (logarithmic scale), speedup, and modularity of communities obtained with GVE-Louvain (Sahu, 2023a) and GVE-LPA for each graph in the dataset.
5.2. Comparing Performance of GVE-LPA
We now compare the performance of GVE-LPA with igraph LPA (Csardi et al., 2006), FLPA (Traag and Šubelj, 2023), and NetworKit LPA (Staudt et al., 2016). igraph LPA and FLPA are both sequential implementations, while NetworKit LPA is parallel. For FLPA, we checkout the suitable branch with modified igraph_community_label_propagation() function, update the label propagation example in C to load the input graph from a file and measure runtime of igraph_community_label_propagation () invoked with IGRAPH_LPA_FAST variant, using gettimeofday(). For igraph LPA, we follow a similar process as FLPA, but checkout code from the master branch of igraph. For NetworKit, we employ a Python script to invoke PLP (Parallel Label Propagation), and measure the total time reported with getTiming(). For each graph, we measure the NetworKit LPA runtime five times for averaging. Due to time constraints and long running times of igraph LPA, we ran it only once per graph. Additionally, we record the modularity of communities obtained from each implementation.
Figure 4(a) shows the runtimes of FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph in the dataset, while Figure 4(b) shows the speedup of GVE-LPA with respect to each aforementioned implementation. igraph LPA fails to run on protein k-mer graphs, i.e., kmer_A2a and kmer_V1r, and hence its results on these graphs are not shown. GVE-LPA exhibits an average speedup of 139×139$\times$139 ×, 97000×97000$\times$97000 ×, and 40×40$\times$40 × over FLPA, igraph LPA, and NetworKit LPA respectively. While FLPA and igraph LPA are sequential implementations, GVE-LPA, running on $64$ threads, is more than 64×64$\times$64 × faster than these implementations. On the sk-2005 graph, GVE-LPA identifies communities in $2.7$ seconds, achieving a processing rate of $1.4$ billion edges/s. Figure 4(c) shows the modularity of communities obtained by each implementation. On average, GVE-LPA obtains $6.6\%$ / $0.2\%$ higher modularity than FLPA (especially on road networks) and igraph LPA respectively, and $4.1\%$ lower modularity than NetworKit LPA (especially on protein k-mer graphs).
Figure 6. Runtime $/|E|$ factor with GVE-LPA for each graph in the dataset.
Figure 7. Average speedup of GVE-Louvain and GVE-LPA with increasing number of threads (in multiples of 2).
Next, we proceed to compare the performance of GVE-LPA with GVE-Louvain (our parallel implementation of Louvain algorithm) in Figure 5. Figures 5(a), 5(b), and 5(c) present the runtimes, speedup (of GVE-LPA with respect to GVE-Louvain), and modularity of GVE-Louvain and GVE-LPA for each graph in the dataset. GVE-LPA offers a mean speedup of 5.4×5.4$\times$5.4 × compared to GVE-Louvain, particularly on social networks, road networks, and protein k-mer graphs, while achieving on average $10.9\%$ lower modularity, especially on social networks and protein k-mer graphs.
In Figure 6, we assess the runtime per edge of GVE-LPA. Results indicate that GVE-LPA exhibits a higher $\text{runtime}/|E|$ factor on graphs with a low average degree (road networks and protein k-mer graphs) and on graphs with a weakly connected clusters (com-LiveJournal and com-Orkut graphs).
5.3. Strong Scaling of GVE-LPA
Finally, we assess the strong scaling performance of GVE-LPA. To this end, we vary the number of threads from $1$ to $64$ in multiples of $2$ for each input graph, and measure the time required for GVE-LPA to identify communities (averaged over five runs). Figure 7 shows the average strong-scaling of of GVE-LPA, in comparison with GVE-Louvain. With 32 threads, GVE-LPA achieves a speedup of 13.5×13.5$\times$13.5 × compared to single-threaded execution, indicating a performance increase of 1.7×1.7$\times$1.7 × for each doubling of threads. At 64 threads, GVE-LPA is moderately impacted by NUMA effects, and offers a speedup of 18.1×18.1$\times$18.1 ×. This contrasts with GVE-Louvain, which appears to be significantly affected by NUMA effects.
6. Conclusion
In summary, this study focuses on optimizing the Label Propagation Algorithm (LPA), a high-speed community detection algorithm, in the shared memory setting. We consider 6 different optimizations, which significantly improve the performance of the algorithm. Comparative assessments against competitive open-source implementations (FLPA) and packages (igraph and NetworKit) indicate that GVE-LPA outperforms FLPA, igraph LPA, and NetworKit LPA by 139×139$\times$139 ×, 97000×97000$\times$97000 ×, and 40×40$\times$40 × respectively, while identifying communities of $6.6\%$ / $0.2\%$ higher quality than FLPA and igraph LPA respectively, and of $4.1\%$ lower quality than NetworKit LPA.
On a web graph with $3.8$ billion edges, GVE-LPA identifies communities in $2.7$ seconds, and thus achieves a processing rate of $1.4$ billion edges/s. GVE-LPA is thus ideal for applications requiring a high-speed clustering algorithm while accepting a compromise in clustering quality — it is on average 5.4×5.4$\times$5.4 × faster than GVE-Louvain, and has a strong scaling factor of 1.7×1.7$\times$1.7 × for every doubling of threads, but identifies communities with on average $10.9\%$ lower modularity than GVE-Louvain.
Acknowledgements.
I would like to thank Prof. Kishore Kothapalli, Prof. Dip Sankar Banerjee, Balavarun Pedapudi, Souvik Karfa, and Vincent Traag for their support.